# Solving Differential Equations using Physics-Informed Deep Equilibrium Models

### v1

@@ Abstract
This paper introduces Physics-Informed Deep Equilibrium Models (PIDEQs) for solving initial value problems (IVPs) of ordinary differential equations (ODEs). Leveraging recent advancements in deep equilibrium models (DEQs) and physics-informed neural networks (PINNs), PIDEQs combine the implicit output representation of DEQs with physics-informed training techniques. We validate PIDEQs using the Van der Pol oscillator as a benchmark problem, demonstrating their efficiency and effectiveness in solving IVPs. Our analysis includes key hyperparameter considerations for optimizing PIDEQ performance. By bridging deep learning and physics-based modeling, this work advances computational techniques for solving IVPs, with implications for scientific computing and engineering applications.
@@ I INTRODUCTION
The advent of deep learning has revolutionized various industries, demonstrating its prowess in tackling complex problems across domains ranging from image recognition to natural language processing. Despite its success, applying deep learning to solve initial value problems (IVPs) of differential equations remains challenging due to the data-driven nature of deep learning models. Gathering sufficient data from real-life dynamical systems can be prohibitively expensive, necessitating a novel approach to training deep learning models for such tasks.
The work by [1] introduced the approach of solving IVPs using deep learning models by optimizing a model’s dynamics rather than solely its outputs. This approach allows deep learning models to approximate the dynamics of a system, provided an accurate description of these dynamics is known, typically in the form of differential equations. By focusing on optimizing the dynamics, the model can be trained with minimal data, covering only the initial and boundary conditions.
@@
This paper explores this integration by studying, implementing, and validating PIDEQs as efficient and accurate solvers for IVPs. An efficient solver is characterized by its ability to operate with minimal data and computational resources, while an effective solver can provide accurate solutions across a wide domain of the independent variable. Our objectives include implementing a DEQ, designing and implementing a PINN training algorithm suitable for DEQs, and evaluating the performance of the physics-informed DEQ in solving IVPs. More specifically, our contributions are:
@@ II-B Physics Regularization
with two distinct components weighted by a scalar parameter $\lambda\in\mathbb{R}^{+}$. By optimizing this cost function, the model is guided to simultaneously satisfy the initial condition and approximate the dynamics represented by $\mathcal{N}$.
@@ III-B Forward Computation
The simplest way to solve the equilibrium equation that defines the output of a DEQ is by iteratively applying the equilibrium function until convergence. Although intuitive, simple iteration is limited as it only works if the equilibrium function is a contraction between the starting point and the equilibrium [4].
Newton’s method offers a faster and more robust alternative by leveraging the Jacobian of $f_{\bm{\theta}}$. As the deep learning paradigm is to perform gradient-based optimization, the well-posedness of the Jacobian of the equilibrium function is already guaranteed. In other words, we can benefit from the fast convergence rate and broad applicability of Newton’s method. In fact, most modern root-finding algorithms can be used for computing the equilibrium, such as Anderson acceleration [5] and Broyden’s Method [6].
@@ V CONCLUSIONS
Future research should focus on evaluating PIDEQs with more complex problems (e.g., higher-order ODEs, PDEs) to better understand their potential advantages. Additionally, exploring more sophisticated methods for the backward pass, such as implicit differentiation techniques, could enhance the training efficiency and accuracy of PIDEQs.
In conclusion, while PIDEQs present a promising direction for integrating physics-informed principles with advanced deep learning architectures, further work is still necessary to fully realize their potential and address their current limitations compared to traditional PINNs. This future research could pave the way for more robust and versatile models capable of solving a wide range of complex dynamical systems, providing competitive solver alternatives.
@@ Acknowledgment
This research was funded in part by Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina (FAPESC) under grant 2021TR2265 and CNPq under grant 402099/2023-0.

### v2

@@ Abstract
This paper introduces Physics-Informed Deep Equilibrium Models (PIDEQs) for solving initial value problems (IVPs) of ordinary differential equations (ODEs). Leveraging recent advancements in deep equilibrium models (DEQs) and physics-informed neural networks (PINNs), PIDEQs combine the implicit output representation of DEQs with physics-informed training techniques. Our validation of PIDEQs, using the Van der Pol oscillator as a benchmark problem, yielded compelling results, demonstrating their efficiency and effectiveness in solving IVPs. Our analysis includes key hyperparameter considerations for optimizing PIDEQ performance. By bridging deep learning and physics-based modeling, this work advances computational techniques for solving IVPs with implications for scientific computing and engineering applications.
@@ I INTRODUCTION
The advent of deep learning has revolutionized various industries, demonstrating its prowess in tackling complex problems across domains ranging from image recognition to natural language processing. Despite the success of deep learning in various industries, applying this technology to solve initial value problems (IVPs) of differential equations presents a formidable challenge, primarily due to the data-driven nature of deep learning models. Gathering sufficient data from real-life dynamical systems can be prohibitively expensive, necessitating a novel approach to training deep learning models for such tasks.
The work by [1] introduced the approach of solving IVPs using deep learning models by optimizing a model’s dynamics rather than solely its outputs. This approach allows deep learning models to approximate the dynamics of a system, provided an accurate description of these dynamics is known, typically in the form of differential equations. Focusing on optimizing the dynamics allows the model to be trained with minimal data, covering only the initial and boundary conditions.
@@
This paper explores this integration by studying, implementing, and validating PIDEQs as efficient and accurate solvers for IVPs. An efficient solver is characterized by its ability to operate with minimal data and computational resources. In contrast, an effective solver can provide accurate solutions across a vast domain of the independent variable. Our research is guided by three objectives: implementing a DEQ, designing and implementing a PINN training algorithm suitable for DEQs, and evaluating the performance of the physics-informed DEQ in solving IVPs. These objectives form the backbone of our study. More specifically, our contributions are:
@@ II-B Physics Regularization
with two distinct components weighted by a scalar parameter $\lambda\in\mathbb{R}_{+}$. By optimizing this cost function, the model is guided to simultaneously satisfy the initial condition and approximate the dynamics represented by $\mathcal{N}$.
@@ III-B Forward Computation
The simplest way to solve the equilibrium equation that defines the output of a DEQ is by iteratively applying the equilibrium function until convergence. Given an input $\bm{x}$ and an initial guess $\bm{z}^{[0]}$, the procedure updates the equilibrium guess $\bm{z}^{[i]}$ by
$\bm{z}^{[i]}=\bm{f}_{\bm{\theta}}(\bm{x},\bm{z}^{[i-1]})$
until $|\bm{z}^{[i]}-\bm{z}^{[i-1]}|$ is sufficiently small. This approach, known as the simple iteration method, is intuitive but can be slow to converge and is sensitive to the starting point. Moreover, it only finds equilibrium of functions that are a contraction between the starting point and the equilibrium point [4]. For example, the simple iteration method fails to find the equilibrium of $f(z)=2z-1$ starting from any point other than the equilibrium itself ($z=1$).
To address these limitations, Newton’s method can be used. It updates the incumbent equilibrium point by
$\bm{z}^{[i+1]}=\bm{z}^{[i]}-\left(\frac{d\bm{f}_{\bm{\theta}}(\bm{x},\bm{z}^{[i]})}{d\bm{z}}\right)^{-1}\bm{f}_{\bm{\theta}}(\bm{x},\bm{z}^{[i]}),$
where $\frac{d}{d\bm{z}}\bm{f}_{\bm{\theta}}(\bm{x},\bm{z}^{[i]})$ represents the Jacobian of $\bm{f}_{\bm{\theta}}$ with respect to $\bm{z}$ [^2] To avoid the computational burden of inverting the Jacobian matrix, it is common to solve $\frac{d\bm{f}_{\bm{\theta}}(\bm{x},\bm{z}^{[i]})}{d\bm{z}}\left(\bm{z}^{[i+1]}-\bm{z}^{[i]}\right)=-\bm{f}_{\bm{\theta}}(\bm{x},\bm{z}^{[i]})$ instead..
Newton’s method converges much faster and is applicable to a broader class of functions compared to simple iteration [4]. It offers a faster and more robust alternative by leveraging the Jacobian of $f_{\bm{\theta}}$. As the deep learning paradigm is to perform gradient-based optimization, the well-posedness of the Jacobian of the equilibrium function is already guaranteed. In other words, we can benefit from the fast convergence rate and broad applicability of Newton’s method. In fact, most modern root-finding algorithms can be used for computing the equilibrium, such as Anderson acceleration [5] and Broyden’s Method [6].
@@ V CONCLUSIONS
Future research should focus on evaluating PIDEQs with more complex problems (e.g., higher-order ODEs and PDEs) to better understand their potential advantages. Additionally, exploring more sophisticated methods for the backward pass, such as implicit differentiation techniques, could enhance the training efficiency and accuracy of PIDEQs.
In conclusion, while PIDEQs present a promising direction for integrating physics-informed principles with advanced deep-learning architectures, further work is still necessary to fully realize their potential and address their current limitations compared to traditional PINNs. This future research could pave the way for more robust and versatile models capable of solving a wide range of complex dynamical systems, providing competitive solver alternatives.
@@ Acknowledgment
This research was funded in part by Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina (FAPESC) under grant 2021TR2265, CNPq under grants 308624/2021-1 and 402099/2023-0, and CAPES under grant PROEX.